# Bayesian Controlled FDR Variable Selection via Knockoffs

### v2

@@ 3.5 Sensitivity analysis
Table 3 shows the results of the sensitivity analysis. We started from the same setting adopted in the simulation study for Scenario 2 and then varied one hyperparameter at a time. We found that posterior inference is robust to the choice of the hyperparameters, except for $v_{0}$ in the prior on the precision matrix $\bm{\Omega}$ in Equation (10). Both Wang, (2015) and Peterson et al., (2016) report the same sensitivity. We notice, however, that the choice of $v_{0}$ does not affect the selection performances of our method. As noted in Barber et al., (2020), knockoffs can still be used with estimated distributions, but the control on FDR is affected by the Kullback-Leiber distance between the estimate and the true distribution. As a matter of fact, in Gaussian graphical models, imposing too many zeros strongly affects the estimates of the non-zero off-diagonal elements on the precision matrix $\bm{\Omega}$. For this reason, we recommend setting the hyperparameters $v_{0}$ and $v_{1}$ to obtain a slightly denser graph.
@@ 4 Application to prostate cancer data
As an illustrative example, we examine a data set on prostate cancer obtained from Stamey et al., (1989) and widely used as a benchmark data set (see, for instance, Tibshirani,, 1996; Zou and Hastie,, 2005, among others). The aim is to investigate the association between Prostate-Specific Antigen (PSA) levels and various clinical measures in a sample of $97$ men, who were preparing to undergo radical prostatectomy. PSA is a protein produced by normal and malignant prostate cells and is useful as a preoperative marker as prostate cancer releases PSA into the bloodstream. Additionally, eight clinical measures are included: the logarithm of the cancer volume (in cm3) (lcavol), the logarithm of the prostate weight (in g) (lweight), patient age (age), the logarithm of the bad prostatic hyperplasia (in cm2) (lbph), the logarithm of capsular penetration (in cm) (lcp), seminal vesicle invasion (svi), the current Gleason score (gleason) and the percentage of Gleason scores four or five out of five (pgg45). The logarithm of the PSA (ng/ml) is the response variable. The presence of seminal vesicle invasion is a binary variable (1 = yes, 0 = no) and gleason is a discrete numerical variable with four values. The Gleason score relates to the current grade of prostate cancer, while the predictor pgg45 provides information about the patient’s history of the Gleason scores, and it is strongly linked to the current final score Gleason score.
@@
For our analysis, we considered the subset of $73$ subjects without seminal vesicle invasion and with no severe cancer (Gleason score less than $9$). To ensure the appropriate statistical assumptions, we applied the nonparanormal transformation (Liu et al.,, 2009) to the covariates. This transformation provides a Gaussian marginal distribution of each covariate and assumes a Gaussian copula on their joint distribution. We tested the joint Gaussian assumption after this transformation using the multivariate Shapiro test (Villasenor Alva and Estrada,, 2009), obtaining a $p$-value of $0.02$. In our application, we used the same hyperparameter settings as in the simulation study, and ran an MCMC chain with $8,000$ iteration for burn-in followed by $8,000$ iterations for inference. The average $z$-scores from the Geweke’s test for $\{(\beta_{j},\widetilde{\beta}_{j})\}_{j=1}^{p}$ was $-0.22$.
@@
Figure 3: Prostate cancer data: Posterior distribution of $W_{j}$, $j=1,\dots,6$. Those highlighted in blue are selected according to Figure 4
@@
Figure 3 shows the posterior distribution of the feature statistics $W_{j}$, $j=1,\ldots,p$ for each covariate in the model, with the selected variables shown in blue. Notably, the posterior distribution of the non-selected variables is symmetric around zero, while the selected variables exhibit a positive-centered distribution. Additionally, Figure 4 reports the estimated values of $2\,\widehat{\mathbb{P}}[W_{j}\leq 0\mid\bm{D}]$, $j=1,\ldots 6$, in increasing order, together with the corresponding estimated $\text{BFDR}(\mathcal{S})$. Therefore, setting $q=0.1$, the covariates to be selected are the logarithm of cancer volume and the logarithm of prostate weight.
@@
We also applied the state-of-art procedures used for comparison in the simulation study. Table 4 summarizes the results. We can see that classical knockoff filter does not find any association between covariates and PSA production level. Spike-and-slab regression selects the same covariates as the proposed method, and the joint estimation model selects one additional variable. Lasso, as we saw in the simulation study, selects many variables, including the two selected by our method. We acknowledge that the Bayesian procedures are computationally more expensive than the classical methods. For example, running the MCMC for our proposed method on this application took 35 minutes on a regular laptop equipped with Intel(R) Core(TM) i7-9750H CPU @ 2.60GHz. However, relying on a posterior distribution of the knockoff variables increases stability of the selection, as our simulations have shown.
@@
Variable BayesKnock Knockoff Filter Lasso Spike-and-slab Joint
@@ 5 Concluding remarks
In this article, we have proposed a fully Bayesian generalization of the model-X knockoff filter of Candès et al., (2018) that allows for variable selection while controlling the BFDR at an arbitrary level. Our framework treats knockoff variables as latent variables that act as negative controls. Additionally, we have imposed a sparse conditional independence structure on the covariates and used it in the prior on the knockoffs in a way that satisfies their pairwise conditional independence property. We have also limited the computational cost by adopting a modified version of the spike-and-slab prior that avoids the increase of the model dimension. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We have shown how our model construction leads to valid model-X knockoffs and demonstrated that the proposed approach is sufficient for controlling the BFDR at an arbitrary level, in finite samples, if the distribution of the covariates is fully known, and asymptotically if estimated as in our model.
@@
In applications, we have shown that our proposal increases the stability of the selection, since it relies on the entire posterior distribution of the knockoff variables instead of a single sample. Results from our simulation study have shown how in simple settings (like Scenarios 1 and 2) our Bayesian knockoff filter approach has similar performance to state-of-art Bayesian variable selection techniques, such as spike-and-slab priors, while outperforming classical variable selection methods, including classical knockoffs filter. Moreover, under more complex dependence structure of the covariates (as in Scenario 3), our proposal is able to control the FDR under any desired threshold, unlike spike-and-slab variable selection methods, that suffer from poor estimates of the PPIs. It also performs as well as classical knockoff filter with covariance structure known. This is relevant since in practice the precision matrix is never known.
@@
Future work could extend this research in several directions. First, while we have assumed a standard Gaussian linear regression setting, similar Bayesian knockoff filters could be designed for linear settings that incorporate discrete-type responses. For example, in the simulations we have also considered an AFT model for survival outcomes. A similar extension, that uses data augmentation techniques, can be used for multinomial probit models (Albert and Chib,, 1993; Sha et al.,, 2004). Furthermore, logit and negative binomial responses could be accommodated via the use of Polya-Gamma data augmentation schemes (Polson et al.,, 2013; Zhou et al.,, 2012). Another interesting direction would be to extend the proposed knockoffs framework to the case of non-Gaussian covariates. The generation of knockoffs for non Gaussian random variables is still an open research problem. This case presents the main challenge of not having a known distribution that respects pairwise exchangeability (Bates et al.,, 2021). A promising direction is represented by the approach proposed by Dreassi et al., (2024), which makes use of conditional independence to build the knockoffs variables. Finally, the variable selection prior distribution we have used on the coefficients presents some computational challenges due to the need to use a Metropolis-within-Gibbs algorithm. Since the largest computational burden comes from the generation of the latent knockoffs at each MCMC iteration, a possible solution would be to use variational inference (Blei et al.,, 2017). However, the restrictive assumptions to obtain valid knockoffs require careful design of the approximate variational distributions. In particular, the commonly used mean field approximation cannot be used since it breaks the pairwise exchangeability assumption of the knockoffs and the covariates.

### v3

@@ 3.4 Sensitivity analysis
Table 3 shows the results of the sensitivity analysis. We started from the same setting adopted in the simulation study for Scenario 2 and then varied one hyperparameter at a time. We found that posterior inference is robust to the choice of the hyperparameters, except for $v_{0}$ in the prior on the precision matrix $\bm{\Omega}$ in Equation (10). Both Wang (2015) and Peterson et al. (2016) report the same sensitivity. We notice, however, that the choice of $v_{0}$ does not affect the selection performances of our method. As noted in Barber et al. (2020), knockoffs can still be used with estimated distributions, but the control on FDR is affected by the Kullback-Leiber distance between the estimate and the true distribution. As a matter of fact, in Gaussian graphical models, imposing too many zeros strongly affects the estimates of the non-zero off-diagonal elements on the precision matrix $\bm{\Omega}$. For this reason, we recommend setting the hyperparameters $v_{0}$ and $v_{1}$ to obtain a slightly denser graph.
@@ 4 Application to prostate cancer data
As an illustrative example, we examine a data set on prostate cancer obtained from Stamey et al. (1989) and widely used as a benchmark data set (see, for instance, Tibshirani, 1996; Zou and Hastie, 2005, among others). The aim is to investigate the association between Prostate-Specific Antigen (PSA) levels and various clinical measures in a sample of $97$ men, who were preparing to undergo radical prostatectomy. PSA is a protein produced by normal and malignant prostate cells and is useful as a preoperative marker as prostate cancer releases PSA into the bloodstream. Additionally, eight clinical measures are included: the logarithm of the cancer volume (in cm3) (lcavol), the logarithm of the prostate weight (in g) (lweight), patient age (age), The logarithm of the amount of benign prostatic hyperplasia (in cm2) (lbph), the logarithm of capsular penetration (in cm) (lcp), seminal vesicle invasion (svi), the current Gleason score (gleason) and the percentage of Gleason scores four or five out of five (pgg45). The logarithm of the PSA (ng/ml) is the response variable. The presence of seminal vesicle invasion is a binary variable (1 = yes, 0 = no) and gleason is a discrete numerical variable with four values. The Gleason score relates to the current grade of prostate cancer, while the predictor pgg45 provides information about the patient’s history of the Gleason scores, and it is strongly linked to the current final score Gleason score.
@@
For our analysis, we considered the subset of $73$ subjects without seminal vesicle invasion and with no severe cancer (Gleason score less than $9$). To ensure the appropriate statistical assumptions, we applied the nonparanormal transformation (Liu et al., 2009) to the covariates. This transformation provides a Gaussian marginal distribution of each covariate and assumes a Gaussian copula on their joint distribution. We tested the joint Gaussian assumption after this transformation using the multivariate Shapiro test (Villasenor Alva and Estrada, 2009), obtaining a $p$-value of $0.02$. In our application, we used the same hyperparameter settings as in the simulation study, and ran an MCMC chain with $8,000$ iteration for burn-in followed by $8,000$ iterations for inference. The average $z$-scores from the Geweke’s test for $\{(\beta_{j},\widetilde{\beta}_{j})\}_{j=1}^{p}$ was 0.22-0.22.
@@
Figure 3: Prostate cancer data: Posterior distribution of $W_{j}$, $j=1,\dots,6$. Those highlighted in black are selected according to Figure 4
@@
Figure 3 shows the posterior distribution of the feature statistics $W_{j}$, $j=1,\ldots,p$ for each covariate in the model, with the selected variables shown in black. Notably, the posterior distribution of the non-selected variables is symmetric around zero, while the selected variables exhibit a positive-centered distribution. Additionally, Figure 4 reports the estimated values of $2\,\widehat{\mathbb{P}}[W_{j}\leq 0\mid\bm{D}]$, $j=1,\ldots 6$, in increasing order, together with the corresponding estimated $\text{BFDR}(\mathcal{S})$. Therefore, setting $q=0.1$, the covariates to be selected are the logarithm of cancer volume and the logarithm of prostate weight.
@@
We also applied the state-of-art procedures used for comparison in the simulation study. Table 4 summarizes the results. We can see that classical knockoff filter does not find any association between covariates and PSA production level. Classical and non-local spike-and-slab regressions select the same covariates as the proposed method, and the joint estimation model selects one additional variable. Lasso, as we saw in the simulation study, selects many variables, including the two selected by our method. We acknowledge that the Bayesian procedures are computationally more expensive than the classical methods. For example, running the MCMC for our proposed method on this application took 35 minutes on a regular laptop equipped with Intel(R) Core(TM) i7-9750H CPU @ 2.60GHz. However, relying on a posterior distribution of the latent knockoff variables increases stability of the selection, as our simulations have shown.
@@
Variable BayesKnock Knockoff Filter Lasso Spike-and-slab Joint Non-local
@@ 5 Concluding remarks
In this article, we have proposed a fully Bayesian generalization of the model-X knockoff filter of Candès et al. (2018) that allows for variable selection while controlling the BFDR at an arbitrary level. Our framework introduces a latent knockoff layer through a parameter-expanded representation of the response model, where knockoffs act as matched negative controls for feature scoring. Additionally, we have imposed a sparse conditional independence structure on the covariates and used it in the prior on the latent knockoffs in a way that satisfies their pairwise conditional independence property. We have also limited the computational cost by adopting a modified version of the spike-and-slab prior that avoids the increase of the model dimension. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We have shown that the latent knockoff variables generated by our construction satisfy the defining properties of valid model-X knockoffs a priori and that the resulting variable-selection procedure controls the BFDR at an arbitrary level in finite samples when the covariate distribution is fully known. When the precision matrix is estimated under the proposed graphical model, the procedure still retains an asymptotic FDR guarantee.
@@
In applications, we have shown that our proposal increases the stability of the selection, since it relies on posterior averaging over the latent knockoff layer rather than on a single realized knockoff sample. Results from our simulation study have shown how in simple settings (like Scenarios 1 and 2) our Bayesian knockoff filter approach has similar performance to state-of-art Bayesian variable selection techniques, such as spike-and-slab priors, while outperforming classical variable selection methods, including classical knockoffs filter. Moreover, under more complex dependence structure of the covariates (as in Scenario 3), our proposal is able to control the FDR under any desired threshold, unlike spike-and-slab variable selection methods, that suffer from poor estimates of the PPIs. It also performs as well as classical knockoff filter with covariance structure known. This is relevant since in practice the precision matrix is never known.
@@
Future work could extend this research in several directions. First, going beyond the standard Gaussian linear regression setting, we have also discussed how similar Bayesian knockoff filters can be designed for linear settings that incorporate discrete-type responses, such as AFT models for survival outcomes and probit models for multinomial responses. More general extensions of the proposed Bayesian knockoff filter formulation might be possible in modeling frameworks that employ latent variables formulations. Another interesting direction would be to extend the proposed knockoffs framework to the case of non-Gaussian covariates. The generation of knockoffs for non Gaussian random variables is still an open research problem. This case presents the main challenge of not having a known distribution that respects pairwise exchangeability (Bates et al., 2021). A promising direction is represented by the approach proposed by Dreassi et al. (2024), which makes use of conditional independence to build the knockoffs variables. Finally, the variable selection prior distribution we have used on the coefficients presents some computational challenges due to the need to use a Metropolis-within-Gibbs algorithm. Since the largest computational burden comes from updating the latent knockoff augmentation at each MCMC iteration, a possible solution would be to use variational inference (Blei et al., 2017). However, the restrictive assumptions to obtain valid knockoffs require careful design of the approximate variational distributions. In particular, the commonly used mean field approximation cannot be used since it breaks the pairwise exchangeability assumption of the knockoffs and the covariates.
Software Availability
Code reproducing the results of this paper will be available on GitHub upon acceptance of the paper.
Acknowledgments
We thank the Editor, Associate Editor, and anonymous referees for their thoughtful and constructive feedback, which greatly improved the quality of this manuscript.
This work was performed as part of the first author’s PhD thesis, and while the author was visiting scholar at Rice University. The first and second authors were partially supported by the Italian Ministry of University and Research (MUR), Department of Excellence project 2023-2027 ReDS ’Rethinking Data Science’ - Department of Statistics, Computer Science, Applications - University of Florence, and the MUR-PRIN grant 2022 SMNNKY, CUP B53D23009470006 European Union - NextGenerationEU, Mission 4 Component, 2. The second author was also partially supported by the European Union - NextGenerationEU - National Recovery and Resilience Plan, Mission 4 Component 2 - Investment 1.5 - THE - Tuscany Health Ecosystem - ECS00000017 - CUP B83C22003920001.